OpenStudy (anonymous):

Use complete sentences to describe the steps taken to simplify this problem. Make sure you include information about the new common denominator, the final simplified expression, and any restrictions. 3/x + 3/x+1 + 3/x+2

OpenStudy (anonymous):

Can you confirm that this is the actual probem\[\frac{3}{x}+\frac{3}{x+1}+\frac{3}{x+2}\]

OpenStudy (anonymous):

Wait sorry i typed it wrong it's \[\frac{ 3 }{ x } - \frac{ 3 }{ x+1 } - \frac{ 3 }{ x+2 }\] sorry : [

OpenStudy (anonymous):

ok, thats good. So now first you multiply the ones in the denominator. So you would get\[x(x+1)(x+2)\]because you need to multiply them all. And then you do the usual subtraction which would ultimately yield this:\[-\frac{3(x^2-2)}{x(x+1)(x+2)}\]The restrictions are that x cannot be 0, -1 or -2.

OpenStudy (anonymous):

How do you subtract them?

OpenStudy (anonymous):

And how do you multiply them so they have the same denominator?

OpenStudy (anonymous):

You just write them in line like that. Much easier to understand what is happening.

OpenStudy (anonymous):

Like what? Sorry but i'm really not following.